rizza's age is one year more than twice norries's age. eight years ago, the sum of their ages was equal to rizza's age ten years from now. how old are they?
Answer by ankor@dixie-net.com(12695)

Let r = R's age now
Let n = N's age now
:
Write an equation for each statement;
:
"rizza's age is one year more than twice norries's age."
r = 2n + 1
:
" eight years ago, the sum of their ages was equal to rizza's age ten years from now."
(r-8)+(n-8) = r + 10
:
r + n - 16 = r + 10
:
r - r + n = 10 + 16
:
n = 26 yrs is N's age now
":
we know r = 2n + 1
r = 2(26) + 1
r = 53 yrs is R's age
:
:
See how those solution work in the statement
"eight years ago, the sum of their ages was equal to rizza's age ten years from now."
(53-8) + (26-8) = 53 + 10
45 + 18 = 63; confirms our solutions

how old are they?